# Code Generation and Algorithmic Problem Solving Using Llama 3.1 405B

### v1

@@ 2 Related Work
LLMs for Related Tasks: With the recent popularity of LLMs, they have been applied for text summarization as well. A recent work (Ahmed and Devanbu, 2023) on code summarization using LLMs used zero-shot and few-shot capabilities of LLMs for project-specific code summarization. A work on book summarization (Chang et al., 2023) tried to develop novel methods to summarize long books using foundational models like LLMs. They divide the entire book into fragments and then follow a hierarchical strategy of divide and conquer to summarize the entire corpus of a long book. LLMs have also been used for comment summarization of videos and topic matching for videos (Dimarco, 2023). Summarization of long meetings has also been attempted using LLMs like GPT-4 and GPT-3.5 and encoder-decoder models like BART (Schneider and Turchi, 2023). There is a work (Maity et al., 2024c) on distractor generation in MCQs where a multi-stage prompting approach delivers better results than a single-stage prompting approach. There are several works on question generation (Maity et al., 2023, 2024b, 2024a; Deroy and Maity, 2024a) for school-level subjects using GPTs. There are several works on classification (Deroy and Maity, 2021; Nigam et al., 2023), summarization (Deroy et al., 2023b, 2024c, a; R. and Sreenivasa Rao, 2022; Ghosh et al., ; Deroy et al., 2021, 2024b; Nigam and Deroy, 2023; Deroy and Maity, 2021; Deroy et al., 2024a), low resource language learning (Maity et al., 2024d), bias detection in texts (Deroy and Maity, 2023), temporal reasoning tasks (Deroy and Maity, 2024b), mathematical reasoning tasks (Deroy and Maity, 2024c) etc. which use GPTs and Gemini-like LLMs.

### v2

@@ 2 Related Work
LLMs and Transformer based models for Related Tasks: With the recent popularity of LLMs, they have been applied for text summarization as well. A recent work (Ahmed and Devanbu, 2023) on code summarization using LLMs used zero-shot and few-shot capabilities of LLMs for project-specific code summarization. A work on book summarization (Chang et al., 2023) tried to develop novel methods to summarize long books using foundational models like LLMs. They divide the entire book into fragments and then follow a hierarchical strategy of divide and conquer to summarize the entire corpus of a long book. LLMs have also been used for comment summarization of videos and topic matching for videos (Dimarco, 2023). Summarization of long meetings has also been attempted using LLMs like GPT-4 and GPT-3.5 and encoder-decoder models like BART (Schneider and Turchi, 2023). There is a work (Maity et al., 2024c) on distractor generation in MCQs where a multi-stage prompting approach delivers better results than a single-stage prompting approach. There are several works on question generation (Maity et al., 2023, 2024b, 2024a; Deroy and Maity, 2024a) for school-level subjects using GPTs. There are several works on classification (Deroy and Maity, 2021; Nigam et al., 2023), summarization (Deroy et al., 2023b, 2024c, a, 2021, 2024b; Nigam and Deroy, 2023; Deroy et al., 2024a), low resource language learning (Maity et al., 2024d), bias detection in texts (Deroy and Maity, 2023), temporal reasoning tasks (Deroy and Maity, 2024b), mathematical reasoning tasks (Deroy and Maity, 2024c) etc. which use GPTs and Gemini-like LLMs and transformer based models.